### v1

# Solving Partial Dominating Set and Related Problems Using Twin-Width

Abstract
Partial vertex cover and partial dominating set are two well-investigated optimization problems. While they are $\rm W[1]$-hard on general graphs, they have been shown to be fixed-parameter tractable on many sparse graph classes, including nowhere-dense classes. In this paper, we demonstrate that these problems are also fixed-parameter tractable with respect to the twin-width of a graph. Indeed, we establish a more general result: every graph property that can be expressed by a logical formula of the form $\phi\equiv\exists x_{1}\ldots\exists x_{k}\#y\,\psi(x_{1},\ldots,x_{k},y)\geq t$, where $\psi$ is a quantifier-free formula, $t$ is an arbitrary number, and #$\#y$ is a counting quantifier, can be evaluated in time $f(d,k)n$, where $n$ is the number of vertices and $d$ is the width of a contraction sequence that is part of the input. Notably, this includes problems such as connected partial dominating set and independent partial dominating set.
keywords:
Partial Dominating Set, Partial Vertex Cover, meta-algorithm, counting logic, twin-width
1 Introduction
Vertex Cover, Independent Set, and Dominating Set are well-known graph problems that have driven significant advancements in parameterized complexity theory. Despite being NP-complete, these problems motivated the development of new approaches, as the difficulty of finding solutions of size $k$ appears to increase significantly from one problem to the next [15]. On general graphs, Vertex Cover has been found to be fixed-parameter tractable, while Independent Set is $\rm W[1]$-complete and Dominating Set is $\rm W[2]$-complete. Consequently, researchers have designed faster algorithms for solving Vertex Cover on general graphs [9, 3, 7, 8, 10, 14, 35, 36, 39], while identifying more and more general graph classes that admit fixed-parameter algorithms for the other two problems. This progression started with planar graphs [1] and led to nowhere-dense graph classes via intermediate results [13, 19, 37, 12]. All three problems can be expressed in first-order logic, and significant advances have been made in meta-algorithm design to solve the first-order model-checking problem. This progression begins with bounded-degree graphs, continues through bounded genus, minor-closed, bounded expansion, nowhere-dense, and finally monadically stable classes [38, 22, 20, 11, 25, 16].
Twin-width is a recently introduced width measure that generalizes clique-width [6]. Classes of bounded twin-width do not contain bounded degree graph classes and are not necessarily monadically stable. They are therefore incomparable to the sparsity hierarchy. Bonnet et al. have developed a parameterized algorithm for first-order model-checking on classes of bounded twin-width [6]. Please note that, unlike tree-width, there is no efficient way known to compute or even approximate twin-width. Moreover, most twin-width based algorithms, including first-order model-checking, require as part of the input not only the graph but also a contraction sequence. The contraction sequence proves that the graph has small twin-width and guides an algorithm. This guidance can be compared to that of dynamic programming on tree decompositions. An optimal tree decomposition can, however, be relatively efficiently computed and even faster approximated. In the world of sparsity algorithms operating in classes of bounded expansion or nowhere-dense classes, the situation is even better: no underlying structure is needed; the algorithm works correctly on every graph but becomes slow when applied outside its scope.
While this represents a significant body of work on Vertex Cover, Independent Set, Dominating Set, and many other first-order expressible problems, the story does not end here. In Vertex Cover, you ask for $k$ nodes that cover all edges, and in Dominating Set, you seek $k$ nodes that dominate all other nodes. Generalizations of these problems are the partial versions: given $k$ and $t$, are there $k$ nodes that cover $t$ edges or are there $k$ nodes that dominate $t$ nodes? These problems are known as Partial Vertex Cover and Partial Dominating Set. As generalizations, they turn out to be harder than their complete counterparts. For example, while Vertex Cover is fixed-parameter tractable on general graphs, Partial Vertex Cover has been shown to be $\rm W[1]$-hard [26], and while Dominating Set is fixed-parameter tractable on degenerate graphs, Partial Dominating Set is $\rm W[1]$-hard [24]. However, it is fixed-parameter tractable with respect to $t$ even on general graphs [29]. Another example of how harder the partial variant can be are $c$-closed graphs, where non-adjacent vertices share at most $c$ neighbors. Dominating Set is fixed-parameter tractable for bounded $c$, while Partial Dominating Set becomes a hard problem even for $c=2$ [28].
Another sequence of papers has demonstrated that Partial Dominating Set can be solved in time $f(k)n^{O(1)}$ for larger and larger graph classes. Amini, Fomin, and Saurabh developed an algorithm with a running time of $O(f(k)\cdot t\cdot n^{C_{H}})$ for the class of $H$-minor-free graphs, where $C_{H}$ is a constant dependent only on $H$. The graph classes with bounded expansion and nowhere-dense graph classes introduced in the sparsity project by Nešetřil and Ossona de Mendez [34] contain all $H$-minor-free classes. For bounded expansion, there exists a meta-algorithm that solves many problems, including Partial Dominating Set in linear fixed-parameter tractable time, i.e., in $f(k)n$ steps [18]. Additionally, for nowhere-dense classes, there is a more restricted meta-algorithm that is general enough to solve Partial Dominating Set in time $O(f(k)n^{1+\epsilon})$ for every $\epsilon>0$ [17].
The “simpler” Partial Vertex Cover has not seen as much attention, but has still been investigated as well; see, for example [31, 21, 32].
Our contribution.
We demonstrate that both Partial Dominating Set and Partial Vertex Cover can be solved efficiently on graph classes with bounded twin-width. To achieve this, we develop two dynamic programming algorithms that work along a contraction sequence provided as part of the input. Interestingly, both algorithms share similarities in their complexity, while usually Partial Vertex Cover is the much easier problem. The algorithms are based on ideas developed by Bonnet et al. in order to solve the (complete) Dominating Set problem on bounded twin-width [5].
Having separate algorithms for related problems is not ideal and it would be beneficial to have them emerge as special cases within a more comprehensive framework. Again, the language of logic could facilitate this. We present a meta-algorithm that solves the model-checking problem for formulas of the form $\phi\equiv\exists x_{1}\cdots\exists x_{k}\#y\psi(x_{1},\ldots,x_{k},y)\geq t$ where $\psi$ must be quantifier-free. Such a formula is true in a graph $G=(V,E)$ if it contains nodes $v_{1},\ldots,v_{k}$ such that $\psi(v_{1},\ldots,v_{k},u)$ holds for at least $t$ many nodes $u\in V$. For instance, with $\psi(x_{1},\ldots,x_{k},y)=\bigvee_{i=1}^{k}(E(x_{i},y)\lor x_{i}=y)$, we can solve Partial Dominating Set. With a quantifier-free $\psi$, we can express numerous other intriguing problems, including Connected Partial Dominating Set and Independent Partial Dominating Set, as well as more exotic ones like “Are there $k$ nodes that are all part of triangles and dominate at least $t$ of all other nodes?” or “How many vertices do we have to delete in order to get a $2$-independent set of size $k$?” (see [27, 2] for applications). The time needed to decide $G\models\phi$ is $d^{O(k^{2}d)}n$, which is linear fpt time. “Linear” means that the input has to be shorter than the classical encoding of the underlying graph as an adjacency list or adjacency matrix. While graphs of bounded twin-width can have a quadratic number of edges, they can still be encoded via their much shorter contraction sequence [23]. Note that our single-purpose algorithm for Partial Dominating Set achieves a better running time of $(kd)^{\mathcal{O}(kd)}n$.
In contrast to Partial Dominating Set, there appears to be no straightforward way to express Partial Vertex Cover within this logic fragment. One common approach in such situations is to subdivide the edges, thereby allowing us to indirectly count edges rather than vertices. This method preserves among others planarity, tree-width, and bounded expansion; however, it does not preserve clique-width or twin-width [4]. The identification of a larger logic framework that enables the efficient solution of more problems and can also express Partial Vertex Cover and related problems remains an open question. Two ways that worked for bounded-expansion and nowhere-dense classes were to use quantifier elimination and logical types. However, quantifier elimination does not work in the context of twin-width and the size of logical types seems to explode.
Twin width has also advantages in comparison to sparsity: If we take graph powers, then its twin-width remains small [6]. This simple observation allows us to tackle the “distance-$r$” variants of Partial Dominating Set as well by using transductions.
2 Preliminaries
For integers $i$ and $j$, we let $[i,j]:=\{n\in\mathbb{N}\colon i\leq n\leq j\}$ and $[i]:=[1,i]$. We study finite, simple and undirected graphs with non-empty vertex set.
Twin-width
A trigraph $G$ is a graph whose edge set is partitioned into a set of black and red edges. The set of red edges is denoted $R(G)$, and the set of black edges $B(G)$; standard graphs are viewed as trigraphs with no red edges. The red graph $G^{R}$ of $G$ is the graph $(V(G),R(G))$ and the black graph $G^{B}$ of $G$ is the graph $(V(G),B(G))$. Any graph-theoretic notion with a color adjective (black or red) has its standard meaning applied in the corresponding graph. For example, the red degree of $u\in V(G)$ in $G$ is its degree in $G^{R}$.
Given a trigraph $G$, a contraction of two distinct vertices $u,v\in V(G)$ is the operation which produces a new trigraph by (1) removing $u,v$ and adding a new vertex $w$, (2) adding a black edge $wx$ for each $x\in V(G)$ such that $xu$, $xv\in B(G)$, and (3) adding a red edge $wy$ for each $y\in V(G)$ such that $yu\in R(G)$, or $yv\in R(G)$, or $y$ contains only a single black edge to either $v$ or $u$. A sequence $C=(G=G_{1},\ldots,G_{n})$ is a contraction sequence of $G$ if it is a sequence of trigraphs such that $|V(G_{n})|=1$ and for all $i\in[n-1]$, $G_{i+1}$ is obtained from $G_{i}$ by contracting two vertices. The width of a contraction sequence $C$ is the maximum red degree over all vertices in all trigraphs in $C$. The twin-width of $G$ is the minimum width of any contraction sequence of $G$. An example of a contraction sequence is provided in Figure 1.
$A$$C$$E$$B$$D$$F$$A$$C$$B$$D$$EF$$AB$$C$$D$$EF$$AB$$CD$$EF$$AB$$CDEF$$ABCDEF$
Figure 1: A contraction sequence of width 2 for the leftmost graph, consisting of $6$ trigraphs.
If a contraction sequence $C=(G_{1},\ldots,G_{n})$ of $G$ is clear from context, then for each $i\in[n]$ and $u\in V(G_{i})$, we call the set of vertices of $G$ contracted to $u$ the bag of $u$ and we denote it $\beta(u)$. Formally, $\beta(u)=\{u\}$ if $u\in V(G)$, and $\beta(u)=\beta(v)\cup\beta(w)$ if there is $i\in[2,n]$ such that $u\in V(G_{i})$, $v,w\in V(G_{i-1})$, and $u$ is created by contracting $v$ and $w$. Note that if a vertex appears in multiple trigraphs in $C$, then its bag is the same in all of them. If $T\subseteq V(G_{i})$ for some $i\in[n]$, then we use $\beta(T)$ to denote $\bigcup_{u\in T}\beta(T)$.
Basic profiles
Now we define basic profiles, inspired by the $k$-profiles in the Dominating Set algorithm from [5]. Intuitively, a profile describes a “type” of a subsolution we compute during our dynamic algorithms. For each profile, we will compute a subsolution that is, in some sense, best for the profile. In the end, we can read the answer to our problem from the profiles associated with the last trigraph $G_{n}$ in the contraction sequence.
In the following sections, we will expand the basic profiles with extra information in two different ways: to extended profiles, which we will use to solve Partial Dominating Set and Partial Vertex Cover in Section 3, and to virtual profiles, which we will use for the model-checking in Section 4. Note that the $k$-profiles in [5] can also be viewed as extensions of our basic profiles. After defining the basic profiles, we will prove some of their properties, which we will later generalize to the extended and virtual profiles.
Let us fix a graph $G$, a contraction sequence $(G=G_{1},\ldots,G_{n})$ of width $d$, and an integer $k\in\mathbb{N}$. For example, in the Partial Dominating Set algorithm, $k$ will be the number of dominating vertices.
Definition 2.1.
If $i\in[n]$, then a basic $i$-profile is a pair $(T,D)$ such that $T\subseteq V(G_{i})$, $|T|\leq k(d+1)$, $T$ is connected in $G_{i}^{R}$, $D\subseteq T$, and $|D|\leq k$.
Figure 2: Right: A basic $i$-profile $\pi=(T,D)$. Vertices of $T$ are colored in gray, and vertices of $D$ have thick boundary. The new vertex of $G_{i}$ is represented by the bigger disc. Left: A $D^{\prime}$-decomposition $\{(T_{1},D_{1}),\ldots,(T_{4},D_{4})\}$ of $\pi$. The sets $T_{1},\ldots,T_{4}$ are represented by the colors blue, green, yellow, and pink. Vertices of $D^{\prime}$ have thick boundary, and the two vertices to be contracted are inside a dashed rectangle.
Now we define how a basic $i$-profile can be decomposed into several basic $(i-1)$-profiles, again inspired by [5]. Later, we will analogously define decompositions for extended and virtual profiles.
Let $i\in[2,n]$, let $v\in V(G_{i})$ be the new vertex in $G_{i}$, let $u_{1},u_{2}\in V(G_{i-1})$ be the two vertices contracted into $v$, let $\pi=(T,D)$ be a basic $i$-profile such that $v\in T$, and let $T^{\prime}=(T\setminus v)\cup\{u_{1},u_{2}\}$. We say that a set $D^{\prime}\subseteq T^{\prime}$ is compatible with $\pi$ if $D\setminus\{v\}=D^{\prime}\setminus\{u_{1},u_{2}\}$, $|D^{\prime}|\leq k$, and $v\in D$ if and only if $u_{1}\in D^{\prime}$ or $u_{2}\in D^{\prime}$. Note that the condition $|D^{\prime}|\leq k$ is necessary only to prohibit the case $|D|=k$ and $u_{1},u_{2}\in D^{\prime}$.
We say that a set $\{(T_{1},D_{1}),\ldots,(T_{m},D_{m})\}$ of basic $(i-1)$-profiles is a $D^{\prime}$-decomposition of $\pi$, see Figure 2, if:
  1. 1.
    The sets $T_{1},\ldots,T_{m}$ are pairwise disjoint.
  2. 2.
    $T^{\prime}=\bigcup_{j\in[m]}T_{j}$, $D^{\prime}=\bigcup_{j\in[m]}D_{j}$, and $D^{\prime}$ is compatible with $\pi$.
  3. 3.
    If $x\in T_{j}$ and $y\in D_{\ell}$ for $j\neq\ell$, then $xy\notin R(G_{i-1})$.
Let us prove that a small decomposition always exists (and can be found efficiently).
Lemma 2.2.
If $i\in[2,n]$, $V(G_{i})\setminus V(G_{i-1})=\{v\}$, $V(G_{i-1})\setminus V(G_{i})=\{u_{1},u_{2}\}$, $\pi=(T,D)$ is a basic $i$-profile such that $v\in T$, $T^{\prime}=(T\setminus v)\cup\{u_{1},u_{2}\}$, and $D^{\prime}\subseteq T^{\prime}$ is compatible with $\pi$, then there is a $D^{\prime}$-decomposition $\mathcal{D}$ of $\pi$ containing at most $d+2$ basic $(i-1)$-profiles. Moreover, $\mathcal{D}$ can be found in time $\mathcal{O}(kd^{2})$.
Proof 2.3.
Let $\{T_{1},\ldots,T_{m}\}$ be the connected components of $T^{\prime}$ in $G_{i-1}^{R}$ and let $D_{j}=T_{j}\cap D^{\prime}$ for each $j\in[m]$. Observe that $(T_{j},D_{j})$ is a basic $(i-1)$-profile if and only if $|T_{j}|\leq k(d+1)$. Hence, if $|T_{j}|\leq k(d+1)$ for each $j\in[m]$, then $\{(T_{1},D_{1}),\ldots,(T_{m},D_{m})\}$ is a $D^{\prime}$-decomposition of $\pi$; condition 3 is satisfied because there are no red edges between distinct $T_{j}$ and $T_{\ell}$. Let $j\in[m]$ and observe that since $T$ is connected in $G_{i}^{R}$, there is a vertex $u\in T_{j}$ such that either $uv\in R(G_{i})$ or $u\in\{u_{1},u_{2}\}$. Since $v$ has red degree at most $d$ in $G_{i}$, we have $m\leq d+2$ as required.
Now suppose that $|T_{j}|>k(d+1)$ for some $j\in[m]$. Since $|T|\leq k(d+1)$ and $T_{j}\subseteq T^{\prime}$, we deduce that $|T_{j}|=k(d+1)+1$, $m=j=1$, and $(T_{j},D_{j})=(T^{\prime},D^{\prime})$. Since each vertex in $D^{\prime}$ has red degree at most $d$ in $G_{i-1}$ and $|D^{\prime}|\leq k$, there is a vertex $w\in T^{\prime}\setminus D^{\prime}$ such that $wx\notin R(G_{i-1})$ for each $x\in D^{\prime}$. Let $\{T_{1}^{\prime},\ldots,T_{p}^{\prime}\}$ be the connected components of $T^{\prime}\setminus\{w\}$ in $G_{i-1}^{R}$ and let $D_{j}^{\prime}=T_{j}^{\prime}\cap D^{\prime}$ for each $j\in[p]$. Now observe that $\{(T_{1}^{\prime},D_{1}^{\prime}),\ldots,(T_{p}^{\prime},D_{p}^{\prime})\}\cup\{(\{w\},\emptyset)\}$ is a $D^{\prime}$-decomposition of $\pi$. Since $T_{j}^{\prime}$ contains a vertex adjacent to $w$ in $G_{i-1}^{R}$ for each $j\in[p]$, we have $p+1\leq d+2$ as required.
Observe that the proof naturally translates into an algorithm finding the desired decomposition and that all the operations can be performed in time linear in $|T^{\prime}|+|R(G_{i-1}[T^{\prime}])|$. Since $|T^{\prime}|\leq k(d+1)+1$ and each vertex of $T^{\prime}$ has red degree at most $d$ in $G_{i-1}$, there are $\mathcal{O}(kd^{2})$ red edges in $G_{i-1}[T^{\prime}]$, and so the stated running time is achieved.
Now we prove that the number of basic $i$-profiles containing any fixed vertex of $G_{i}$ is small.
Lemma 2.4.
If $i\in[n]$ and $v\in V(G_{i})$, then the number of basic $i$-profiles $(T,D)$ such that $v\in T$ is in $d^{\mathcal{O}(kd)}$. Moreover, such profiles can be enumerated in time $d^{\mathcal{O}(kd)}$.
Proof 2.5.
Let $(T,D)$ be a basic $i$-profile such that $v\in T$. By Corollary 3.2. from [5], there are at most $d^{2\ell-2}+1$ red-connected subsets of $V(G_{i})$ of size $\ell$ containing $v$. Since $|T|\leq k(d+1)$, there are $d^{2kd+2k-2}+1$ possible choices of $T$. Since $D\subseteq T$, there are $2^{k(d+1)}$ possible choices of $D$. Hence, there are $d^{\mathcal{O}(kd)}$ basic $i$-profiles containing $v$. The claim about enumeration follows from Corollary 3.2. as well.
3 Partial Dominating Set and Vertex Cover
In this section, we show that the following parameterized problems are fixed-parameter tractable. When a graph $G$ is clear from context, we denote the closed neighborhood of $S\subseteq V(G)$ in a $G$ by $N[S]$, i.e., $N[S]=S\cup\{u\in V(G)\colon u$ has a neighbor in $S\}$.
Partial Dominating Set using twin-width
Input: A graph $G$, a contraction sequence for $G$ of width $d$, integers $k$ and $t$
Question: Is there a set $S\subseteq V(G)$ of size $k$ such that $|N[S]|\geq t$?
Parameter: $k+d$
Partial Vertex Cover using twin-width
Input: A graph $G$, a contraction sequence for $G$ of width $d$, integers $k$ and $t$
Question: Is there a set $S\subseteq V(G)$ of size $k$ such that at least $t$ edges have an endpoint in $S$?
Parameter: $k+d$
Henceforth, we refer to these problems simply as Partial Dominating Set and Partial Vertex Cover.
3.1 Extended Profiles
For both of these problems, we will need a notion of a profile, obtained by extending basic profiles introduced in Section 2. Let us fix a graph $G$, a contraction sequence $(G=G_{1},\ldots,G_{n})$ of width $d$, and an integer $k$.
Definition 3.1.
Let $i\in[n]$ and let $\pi=(T,D)$ be basic $i$-profile. An extended $i$-profile is a tuple $\pi^{\prime}=(T,D,M,f)$ such that $M\subseteq T$ and $f\colon T\rightarrow[0,k]$ is a function such that $\sum_{u\in T}f(u)\leq k$ and for each $u\in T$, $f(u)\leq|\beta(u)|$, and $D=\{u\in T\colon f(u)\neq 0\}$.
Informally, the function $f$ specifies the distribution of the dominating (resp. covering) vertices across $D$, and the set $M$ specifies which bags we want to dominate vertices in. Note that for Partial Vertex Cover, $M$ will not be used, i.e., $M=\emptyset$ will always hold (this allows us to use the same definition of a profile for both problems). The following definition formalizes this idea.
Definition 3.2.
Let $i\in[n]$ and let $\pi=(T,D,M,f)$ be an extended $i$-profile. A set $S\subseteq\beta(T)$ is a solution of $\pi$ if for every $u\in T$, $f(u)=|S\cap\beta(u)|$. We define the dominating-set-value of $S$ in $\pi$ as $|N[S]\cap\beta(M)|$ and the vertex-cover-value of $S$ in $\pi$ as |$|\{uv\in E(G)\colon u\in S,v\in\beta(T)\}|$.
Note that each set $S\subseteq V(G)$ can be a solution of more than one extended $i$-profile for each $i\in[n]$. Indeed, $S$ determines only $D$ and $f$, whereas there are more choices of $T$ and $M$ ($T$ can be any small red-connected set containing $D$). However, it will always be clear from context which profile a solution belongs to, and so we will often simply say, e.g., the dominating-set-value of $S$ (omitting “in $\pi$”).
Let us give a quick informal overview of how the profiles will be used to solve Partial Dominating Set. During the algorithm, we proceed through each time $i$ ranging from $1$ to $n$ and compute for each extended $i$-profile $\pi$ its optimal value; that is, the maximum value $t$ such that $S$ dominates $t$ vertices from $\beta(M)$. To compute the optimal value, we need to decompose the profile: each decomposition is a small set of extended $(i-1)$-profiles, see Section 3.2 for a formal definition. Crucially, each solution of $\pi$ is described by some decomposition, and the best solution described by a decomposition $\mathcal{D}=\{\pi_{1},\ldots,\pi_{m}\}$ can be computed from the optimal solutions of the profiles in $\mathcal{D}$ (which were computed in the previous iteration). Hence, the optimal solution of $\pi$ can be computed by trying all possible decompositions. More precisely, we do not try all decompositions but only one decomposition of each “type”. Finally, to determine how many vertices can be dominated by a set of size $k$ in $G$, it suffices to examine the optimal value of one particular $n$-profile $\pi$ (note that $G_{n}$ consists of a single vertex, and so it is easy to define $\pi$ explicitly). We get this value at the end of the run of Algorithm 1.
Let us now show that the number of extended $i$-profiles containing any fixed vertex of $G_{i}$ is bounded.
Lemma 3.3.
If $i\in[n]$ and $v\in V(G_{i})$, then the number of extended $i$-profiles $(T,D,M,f)$ such that $v\in T$ is in $2^{\mathcal{O}(kd\log(kd))}$. Moreover, such profiles can be enumerated in time $2^{\mathcal{O}(kd\log(kd))}$.
Proof 3.4.
By Lemma 2.4, there is $d^{\mathcal{O}(kd)}$ basic $i$-profiles $(T,D)$. Recall that $|T|\leq k(d+1)$. Since $M\subseteq T$, there are $2^{k(d+1)}$ possible choices of $M$. Since $f$ is a function of type $T\rightarrow[0,k]$, there are at most $(k+1)^{k(d+1)}$ possible choices of $f$. Hence, there are $d^{\mathcal{O}(kd)}\cdot 2^{\mathcal{O}(kd)}\cdot k^{\mathcal{O}(kd)}=2^{\mathcal{O}(kd\log(kd))}$ extended $i$-profiles containing $v$, and the claim about enumeration follows from Lemma 2.4.
3.2 Extended Decompositions
Now we generalize the $D^{\prime}$-decompositions defined in Section 2 to extended profiles.
Let $i\in[2,n]$, let $v\in V(G_{i})$ be the new vertex in $G_{i}$, let $u_{1},u_{2}\in V(G_{i-1})$ be the two vertices contracted into $v$, let $\pi_{0}=(T,D)$ be a basic $i$-profile such that $v\in T$, let $T^{\prime}=T\setminus\{v\}\cup\{u_{1},u_{2}\}$, and let $\pi=(T,D,M,f)$ be an extended $i$-profile. We say that a function $f^{\prime}\colon T^{\prime}\rightarrow[0,k]$ is compatible with $\pi$ if $f(v)=f^{\prime}(u_{1})+f^{\prime}(u_{2})$, $f^{\prime}(u_{1})\leq|\beta(u_{1})|$, $f^{\prime}(u_{2})\leq|\beta(u_{2})|$, and for each $u\in T\setminus\{v\}$, $f(u)=f^{\prime}(u)$. Note that there is at least one function compatible with $\pi$ because $f(v)\leq|\beta(v)|$. Let $M^{*}=M\setminus\{v\}\cup\{u_{1},u_{2}\}$ if $v\in M$, and $M^{*}=M$ otherwise. Observe that $M^{*}\subseteq T^{\prime}\subseteq V(G_{i-1})$ and let $M^{\prime}$ be the set defined as follows.
$M^{\prime}=\{u\in M^{*}\colon u\text{ has no black neighbor in }D^{\prime}\text{ in }G_{i-1}\}.$ (1)
Let $\mathcal{D}=\{(T_{j},D_{j},M_{j},f_{j})\colon j\in[m]\}$ be a set of extended $(i-1)$-profiles such that $M^{\prime}=\bigcup_{j\in[m]}M_{j}$, and let $D^{\prime}=\bigcup_{j\in[m]}D_{j}$ and $f^{\prime}:=\bigcup_{j\in[m]}f_{j}$. Observe that $f^{\prime}$ is a function of type $T^{\prime}\rightarrow[0,k]$ because $T_{1},\ldots,T_{m}$ are pairwise disjoint. We say that $\mathcal{D}$ is an $f^{\prime}$-decomposition of $\pi$ if $\{(T_{1},D_{1}),\ldots,(T_{m},D_{m})\}$ is a $D^{\prime}$-decomposition of $\pi_{0}$ and $f^{\prime}$ is compatible with $\pi$. This terminology makes sense because $D^{\prime}$ is determined by $f^{\prime}$, and $T^{\prime}$ and $M^{\prime}$ are determined by $T$ and $M$, respectively.
Now we prove that a small $f^{\prime}$-decomposition exists for each $f^{\prime}$ compatible with $\pi$.
Lemma 3.5.
If $i\in[2,n]$, $V(G_{i})\setminus V(G_{i-1})=\{v\}$, $V(G_{i-1})\setminus V(G_{i})=\{u_{1},u_{2}\}$, $\pi=(T,D,M,f)$ is an extended $i$-profile such that $v\in T$, $T^{\prime}=T\setminus\{v\}\cup\{u_{1},u_{2}\}$, and $f^{\prime}\colon T^{\prime}\rightarrow[0,k]$ is a function compatible with $\pi$, then there is an $f^{\prime}$-decomposition $\mathcal{D}$ of $\pi$ with cardinality at most $d+2$. Moreover, $\mathcal{D}$ can be found in time $\mathcal{O}(k^{2}d^{2})$.
Proof 3.6.
Let $D^{\prime}=\{u\in T^{\prime}\colon f^{\prime}(u)\neq 0\}$, let $M^{*}=M\setminus\{v\}\cup\{u_{1},u_{2}\}$ if $v\in M$, and $M^{*}=M$ otherwise, and let $M^{\prime}\subseteq T^{\prime}$ be the set satisfying Equation 1; note that $M^{\prime}$ is determined by $D^{\prime}$, which in turn depends on $f^{\prime}$. By Lemma 2.2, there is a $D^{\prime}$-decomposition $\mathcal{D}_{0}=\{(T_{1},D_{1}),\ldots,(T_{m},D_{m})\}$ of the basic $i$-profile $(T,D)$ such that $m\leq d+2$. For each $j\in[m]$, let $M_{j}=M^{\prime}\cap T_{j}$ and let $f_{j}$ be the restriction of $f^{\prime}$ to $T_{j}$. It can be easily observed that $\{(T_{j},D_{j},M_{j},f_{j})\colon j\in[m]\}$ is the desired $f^{\prime}$-decomposition of $\pi$.
By Lemma 2.2, $\mathcal{D}_{0}$ can be found in time $\mathcal{O}(kd^{2})$. Since $|M^{*}|\leq|T^{\prime}|\leq k(d+1)+1$ and $|D^{\prime}|\leq k$, there are $\mathcal{O}(k^{2}d)$ possible black edges one has to inspect when computing $M^{\prime}$. Once $M^{\prime}$ is computed, deriving $M_{j}$ (and $f_{j}$) for each $j\in[m]$ is trivial, which implies the stated running time.
The following lemma shows that a solution of a profile $\pi$ can be constructed by “merging” the solutions of profiles in a decomposition of $\pi$.
Lemma 3.7.
If $i\in[2,n]$, $\pi$ is an extended $i$-profile, $\{\pi_{j}\colon j\in[m]\}$ is an $f^{\prime}$-decomposition of $\pi$, and $S_{j}$ is a solution of $\pi_{j}$ for each $j\in[m]$, then $S:=\bigcup_{j\in[m]}S_{j}$ is a solution of $\pi$.
Proof 3.8.
Let $\pi=(T,D,M,f)$ and for each $j\in[m]$, let $\pi_{j}=(T_{j},D_{j},M_{j},f_{j})$. By Definition 3.2, $S_{j}\subseteq\beta(T_{j})$ for each $j\in[m]$. Since $T_{1},\ldots,T_{m}$ are pairwise disjoint, $S_{1},\ldots,S_{m}$ are pairwise disjoint as well. Observe that $S\subseteq\bigcup_{j\in[m]}\beta(T_{j})=\beta(T)$ as required.
Now we need to show that $f(u)=|S\cap\beta(u)|$ for each $u\in T$. Let $v\in V(G_{i})$ be the new vertex in $G_{i}$ and let $u_{1},u_{2}\in V(G_{i-1})$ be the two vertices contracted into $v$. First, let $u\in T\setminus\{v\}$. Observe that there is $j\in[m]$ such that $u\in T_{j}$ and $f_{j}(u)=f^{\prime}(u)=f(u)$ because $f^{\prime}$ is compatible with $\pi$. Since $S_{j}$ is a solution of $\pi_{j}$, we have $f(u)=|S_{j}\cap\beta(u)|=|S\cap\beta(u)|$ as required. Second, let $j,\ell\in[m]$ be such that $u_{1}\in T_{j}$ and $u_{2}\in T_{\ell}$. Since $f^{\prime}$ is compatible with $\pi$, we have $f(v)=f^{\prime}(u_{1})+f^{\prime}(u_{2})$. Since $S_{j}$ (resp. $S_{\ell}$) is a solution of $\pi_{j}$ (resp. $\pi_{\ell}$), we obtain $f(v)=|S_{j}\cap\beta(u_{1})|+|S_{\ell}\cap\beta(u_{2})|=|S\cap(\beta(u_{1})\cup\beta(u_{2}))|=|S\cap\beta(v)|$ as required.
3.3 Partial Dominating Set
Let $(G,C=(G_{1},\ldots,G_{n}),k,t)$ be an instance of the Partial Dominating Set problem; recall that $C$ is a contraction sequence of width $d$. In this section, we show that Algorithm 1 solves the problem in time $(dk)^{\mathcal{O}(dk)}n$, see Theorem 3.15. For an outline of the algorithm, see Section 3.1.
We will use the notion of extended profiles, see Definition 3.1, and the notion of solutions, see Definition 3.2. Instead of the dominating-set-value of a solution, we will say simply the value. We say that a solution $S$ is an optimal solution of an extended profile $\pi$ if no other solution of $\pi$ has higher value than $S$.
Before presenting the algorithm, let us make two key observations.
Lemma 3.9.
Let $i\in[2,n]$, $\pi$ be an extended $i$-profile, $\{(T_{j},D_{j},M_{j},f_{j})\colon j\in[m]\}$ be an $f^{\prime}$-decomposition of $\pi$, $S$ be a solution of $\pi$, and $S_{j}=S\cap\beta(T_{j})$ for each $j\in[m]$. For each $j\in[m]$:
  1. a)
    The value of $S_{j}$ equals $|N[S]\cap\beta(M_{j})|$.
  2. b)
    If $u\in T_{j}$ has a black neighbor in $D_{\ell}$ for some $\ell\in[m]$, then $\beta(u)\subseteq N[S]$.
Proof 3.10.
For a), let us fix $j\in[m]$. It suffices to prove that $N[S_{j}]\cap\beta(M_{j})=N[S]\cap\beta(M_{j})$. Suppose for contradiction that there are vertices $u_{0}\in\beta(M_{j})$ and $v_{0}\in S\setminus S_{j}$ such that $u_{0}v_{0}\in E(G)$. Let $u,v\in V(G_{i-1})$ be such that $u_{0}\in\beta(u)$ and $v_{0}\in\beta(v)$, and observe that $v\in D_{\ell}$ for some $\ell\in[m]$, $\ell\neq j$. Since $u\in M_{j}$, $uv$ is not a black edge in $G_{i-1}$, see Equation (1). However, $uv$ is not a red edge in $G_{i-1}$ either, by condition 3 in the definition of a basic decomposition in Section 2. This is a contradiction with $u_{0}v_{0}\in E(G)$.
For b), let us fix $j\in[m]$ and $u\in T_{j}$ that has a black neighbor $v\in D_{\ell}$ for some $\ell\in[m]$. By Definition 3.1, $f^{\prime}(v)\neq 0$ and there is a vertex $w\in\beta(v)\cap S$. Hence, $\beta(u)\subseteq N(w)\subseteq N[S]$ as required.
Input: A graph $G$, contraction sequence $(G_{1}=G,\ldots,G_{n})$ of width $d$, and $k,t\in\mathbb{N}$.
1 for each extended $1$-profile $\pi=(T,D,M,f)$ do  $\sigma_{1}(\pi):=(D,|D\cap M|)$
2 for $i$ from $2$ to $n$ do
3       $v:=$ the new vertex in $G_{i}$
4       $u_{1},u_{2}:=$ the two vertices of $G_{i-1}$ contracted into $v$
5       for each extended $i$-profile $\pi=(T,D,M,f)$ do
6             if $v\notin T$ then $\sigma_{i}(\pi):=\sigma_{i-1}(\pi)$; continue with next $\pi$
7             $\sigma_{i}(\pi):=(\textsf{null},-1)$
8             $T^{\prime}:=(T\setminus v)\cup\{u_{1},u_{2}\}$
9             if $v\in M$ then $M^{*}:=M\setminus\{v\}\cup\{u_{1},u_{2}\}$ else $M^{*}:=M$
10             for each function $f^{\prime}\colon T^{\prime}\rightarrow[0,k]$ compatible with $\pi$ do
11                   Let $\{\pi_{j}\colon j\in[m]\}$ be an $f^{\prime}$-decomposition of $\pi$ s.t. $m\leq d+2$; it exists by Lemma 3.5.
12                   for each $j\in[m]$ do
13                         $(T_{j},D_{j},M_{j},f_{j}):=\pi_{j}$; $(S_{j},\nu_{j}):=\sigma_{i-1}(\pi_{j})$
14                         $C_{j}=\{u\in M^{*}\cap T_{j}\colon u$ has a black neighbor in $D_{\ell}$ for some $\ell\in[m]\}$
15                         $\textsf{value}_{j}:=\nu_{j}+\sum_{u\in C_{j}}|\beta(u)|$
16
17                  $\nu:=\sum_{j\in[m]}\textsf{value}_{j}$; $S:=\bigcup_{j\in[m]}S_{j}$
18                   if $\nu\geq\nu_{0}$, where $(S_{0},\nu_{0})=\sigma_{i}(\pi)$ then $\sigma_{i}(\pi):=(S,\nu)$
19
20
21
22$(S,\nu):=\sigma_{n}(\{u\},\{u\},\{u\},\{(u,k)\})$, where $\{u\}:=V(G_{n})$
23 return $\nu\geq t$
Algorithm 1 Algorithm for Partial Dominating Set
Now we show that in each iteration of the second innermost loop of Algorithm 1, a solution of the currently processed profile $\pi$ and its value are computed. Informally, for each profile $\pi_{j}$ in the decomposition of $\pi$, we avoid double-counting dominated vertices in $\beta(M^{*}\cap T_{j})$ by partitioning $M^{*}\cap T_{j}$ into two subsets, $M_{j}$ and $C_{j}$: $\beta(M_{j})$ contains vertices dominated “from inside”, i.e., via an edge considered in one of the previous iterations, and $\beta(C_{j})$ contains those dominated “from outside”, i.e., via a black edge to $D^{\prime}$.
Lemma 3.11.
Let $i\in[2,n]$, $\pi$ be an extended $i$-profile, $f^{\prime}$ be a function compatible with $\pi$, and $\nu$ and $S$ be the values computed by Algorithm 1 on line 1 in the iteration processing $f^{\prime}$. If for each extended $(i-1)$-profile $\pi^{\prime}$, $S^{\prime}$ is a solution of $\pi^{\prime}$ of value $\nu^{\prime}$, where $(S^{\prime},\nu^{\prime})$ is the pair computed in $\sigma_{i-1}(\pi^{\prime})$, then $S$ is a solution of $\pi$ of value $\nu$.
Proof 3.12.
Let $\pi=(T,D,M,f)$ and let $\{\pi_{j}\colon j\in[m]\}$ be the $f^{\prime}$-decomposition of $\pi$ found on line 1 in the iteration processing $f^{\prime}$. For each $j\in[m]$, let $T_{j},D_{j},M_{j},f_{j},S_{j},\nu_{j}$, and $C_{j}$ be the values computed on lines 1 and 1. By Lemma 3.7, $S$ is a solution of $\pi$. To simplify notation, let $\delta(A):=|N[S]\cap\beta(A)|$ for any $A\subseteq V(G_{i})$ or $A\subseteq V(G_{i-1})$. Now it suffices to show that $\nu=\delta(M)$.
Let $M^{*}$ be defined as in the algorithm, see line 1, and observe that $\delta(M)=\delta(M^{*})$. Since $T_{1},\ldots,T_{m}$ are pairwise disjoint, it suffices to show that for each $j\in[m]$, $\textsf{value}_{j}$ computed on line 1 equals $\delta(M^{*}\cap T_{j})$. Let us fix $j\in[m]$, and observe that $M^{*}\cap T_{j}=M_{j}\cup C_{j}$ and $M_{j}\cap C_{j}=\emptyset$, see Equation (1) and line 1. By Lemma 3.9a, we have $\nu_{j}=\delta(M_{j})$ because $\nu_{j}$ is the value of $S_{j}$. By Lemma 3.9b, $\beta(u)\subseteq N[S]$ for each $u\in C_{j}$, which implies $\delta(C_{j})=|\beta(C_{j})|$. Hence, $\textsf{value}_{j}=\delta(M_{j})+\delta(C_{j})=\delta(M^{*}\cap T_{j})$, and $\nu=\delta(M)$ as required.
Now we show that we are able to find an optimal solution for each profile. The idea is to describe an optimal solution with a function $f^{\prime}$ compatible with the considered profile, and then to show that the solution found in the iteration processing $f^{\prime}$ is optimal as well.
Lemma 3.13.
For each $i\in[n]$ and each extended $i$-profile $\pi$, if the final value of $\sigma_{i}(\pi)$ computed by Algorithm 1 is $(S,\nu)$, then $S$ is an optimal solution of $\pi$ and $\nu$ is the value of $S$.
Proof 3.14.
Let $i\in[n]$ and $\pi=(T,D,M,f)$ be an extended $i$-profile. We proceed by induction on $i$. First, suppose that $i=1$. Since there are no red edges in $G_{1}=G$ and $T$ is connected in $G^{R}$, we know that $T=\{u\}$ for some $u\in V(G)$. Notice that $D,M\in\{\emptyset,\{u\}\}$ and $f\in\{\{(u,0)\},\{(u,1)\}\}$. In particular, $\pi$ has only one solution, namely $D$, and the value of $D$ is $|D\cap M|$ as required, see line 1. Now suppose that $i>1$. Let $v\in V(G_{i})$ be the new vertex in $G_{i}$ and let $u_{1},u_{2}\in V(G_{i-1})$ be the two vertices contracted into $v$. If $v\notin T$, then $\pi$ is also an $(i-1)$-profile. By induction hypothesis, $\sigma_{i-1}(\pi)$ is an optimal solution of $\pi$, which ensures that $\sigma_{i}(\pi)$ is also an optimal solution of $\pi$, see line 1. Hence, let us assume that $v\in T$ and let $T^{\prime}=(T\setminus v)\cup\{u_{1},u_{2}\}$. By induction hypothesis and Lemma 3.11, $\pi$ has a solution since there is a function compatible with $\pi$.
Let $S^{*}$ be an optimal solution of $\pi$, let $\nu^{*}$ be the value of $S^{*}$, and let $f^{\prime}\colon T^{\prime}\rightarrow[0,k]$ be defined as $f^{\prime}(u)=|S^{*}\cap\beta(u)|$. Observe that by Definition 3.2, $f^{\prime}(u)=f(u)$ for each $u\in T\setminus\{v\}$, and $f(v)=|S^{*}\cap\beta(v)|=|S^{*}\cap\beta(u_{1})|+|S^{*}\cap\beta(u_{2})|=f^{\prime}(u_{1})+f^{\prime}(u_{2})$, i.e., $f^{\prime}$ is compatible with $\pi$. Let us consider the iteration of the loop on lines 1-1 processing $\pi$ and the iteration of the loop on lines 1-1 processing $f^{\prime}$. Let $\{\pi_{j}\colon j\in[m]\}$ be the $f^{\prime}$-decomposition of $\pi$ found on line 1. For each $j\in[m]$, let $T_{j},D_{j},M_{j},f_{j},S_{j},\nu_{j}$, and $C_{j}$ be as in the algorithm. Let $\nu$ and $S$ be the values computed on line 1. By Lemma 3.11, $S$ is a solution of $\pi$ with value $\nu$. Similarly to the proof of Lemma 3.11, we denote $\delta(A)=|N[S]\cap\beta(A)|$ and $\delta^{*}(A)=|N[S^{*}]\cap\beta(A)|$ for any $A\subseteq V(G_{i})$ or $A\subseteq V(G_{i-1})$.
Suppose for contradiction that $\nu<\nu^{*}$, i.e., $\delta(M)<\delta^{*}(M)$, see Definition 3.2. Let $M^{*}$ be as in the algorithm, see line 1, and observe that $\delta(M)=\delta(M^{*})$ and $\delta^{*}(M)=\delta^{*}(M^{*})$. Hence, there must be $j\in[m]$ such that $\delta(M^{*}\cap T_{j})<\delta^{*}(M^{*}\cap T_{j})$; let us fix such a $j$. By Lemma 3.9b, if $u\in C_{j}$, then $\beta(u)\subseteq N[S]\cap N[S^{*}]$. Hence, $\delta(C_{j})=|\beta(C_{j})|=\delta^{*}(C_{j})$, which implies $\delta(M_{j})<\delta^{*}(M_{j})$ because $M^{*}\cap T_{j}=C_{j}\cup M_{j}$ and $M_{j}\cap C_{j}=\emptyset$. Let $S_{j}^{*}=S^{*}\cap\beta(T_{j})$ and observe that $S_{j}^{*}$ is a solution of $\pi_{j}$. By Lemma 3.9a, $\delta(M_{j})$ is the value of $S_{j}$ and $\delta^{*}(M_{j})$ is the value of $S_{j}^{*}$. By induction hypothesis, $S_{j}$ is an optimal solution of $\pi_{j}$, which is a contradiction with $\delta(M_{j})<\delta^{*}(M_{j})$. Hence, $\nu=\nu^{*}$, and $S$ is an optimal solution of $\pi$, too.
After the iteration of the loop on lines 1-1 processing $f^{\prime}$, $\sigma_{i}(\pi)=(S,\nu)$ because $S$ is an optimal solution of $\pi$. Observe that the second component of $\sigma_{i}(\pi)$ cannot change in any of the following iterations because $\nu$ is always the value of some solution of $\pi$ by Lemma 3.11. Hence, after all functions compatible with $\pi$ have been processed, $\sigma_{i}(\pi)$ is indeed an optimal solution of $\pi$ and its value, as desired.
Using Lemma 3.13, it is easy to show the correctness of Algorithm 1.
Theorem 3.15.
Algorithm 1 decides the Partial Dominating Set problem in time $2^{\mathcal{O}(kd\log(kd))}\cdot n$.
Proof 3.16.
Let $\iota=(G,C=(G=G_{1},\ldots,G_{n}),k,t)$ be the input instance, where $C$ is contraction sequence of width $d$. Let $V(G_{n})=\{u\}$ and let $\pi=(\{u\},\{u\},\{u\},\{(u,k)\})$. Note that $\pi$ is an extended $n$-profile unless $n<k$, in which case, $\iota$ is trivially a NO instance. By Lemma 3.13, $S$ is an optimal solution of $\pi$ with value $\nu$, where $(S,\nu)$ is the pair computed by Algorithm 1 on line 1. By Definition 3.2, $|S|=k$ and $\nu=|N[S]\cap\beta(\{u\})|=|N[S]|$. Hence, $\iota$ is a YES instance if and only if $\nu\geq t$, as required.
Now need to argue that Algorithm 1 achieves the desired running time. As an optimization, we maintain a single function $\sigma$ (instead of $\sigma_{i}$ for each $i\in[n]$). Since there are $\mathcal{O}(n)$ extended $1$-profiles, line 1 can be performed in time $\mathcal{O}(n)$. There are $n-1$ iterations of the outermost loop (i.e., the loop on lines 1-1); let us consider the iteration processing $i\in[2,n]$ and let $v$ be the new vertex in $G_{i}$. Let $\pi=(T,D,M,f)$ be an extended $i$-profile. Since $\sigma(\pi)$ has already been computed if $v\notin T$, it suffices to consider only profiles such that $v\in T$; these profiles can be enumerated in time $2^{\mathcal{O}(kd\log(kd))}$ by Lemma 3.3. There are at most $k+1$ functions $f^{\prime}$ compatible with $\pi$ and for each of them, we can find an $f^{\prime}$-decomposition of size at most $m\leq d+2$ in time $\mathcal{O}(k^{2}d^{2})$ by Lemma 3.5. Hence, the running time of one iteration of the outermost loop is $2^{\mathcal{O}(kd\log(kd))}$, which implies that the total running time is $2^{\mathcal{O}(kd\log(kd))}\cdot n$ as required.
3.4 Partial Vertex Cover
In this section, we want to adapt the previous result to the Partial Vertex Cover problem.
Theorem 3.17.
There is an algorithm that, given a graph $G$, a contraction sequence $(G_{1}=G,\ldots,G_{n})$ of width $d$, and two integers $k$ and $t$, decides the Partial Vertex Cover problem in time $2^{\mathcal{O}(kd\log(kd))}\cdot n$.
The proof of the theorem above is similar to the proof of Theorem 3.15. Hence, we first discuss informally the differences between the two proofs and only afterwards give the formal proof in Section 3.4. The first difference is already mentioned in Section 3.1; it suffices to consider extended $i$-profiles $\pi=(T,D,M,f)$ such that $M=\emptyset$.
Recall that the vertex-cover-value of a solution $S\subseteq\beta(T)$ is the number of edges in $G[\beta(T)]$ covered by $S$, see Definition 3.2. Hence, if $\mathcal{D}=\{(T_{j},D_{j},\emptyset,f_{j})\colon j\in[m]\}$ is a decomposition of $\pi$, then the edges in $G[\beta(T_{j})]$ covered by $S$ are covered also by $S\cap\beta(T_{j})$. This means that when merging solutions of profiles in $\mathcal{D}$, we must only add the number $\delta_{j\ell}$ of covered edges with one endpoint in $\beta(T_{j})$ and the other in $\beta(T_{\ell})$ for distinct $j,\ell\in[m]$. Since there cannot be a red edge connecting $D_{j}$ and $T_{\ell}$ by definition of a basic decomposition, we can compute $\delta_{j\ell}$ by inspecting the black edges between $T_{j}$ and $T_{\ell}$. Observe that a black edge $uv\in E(G_{i-1})$ semi-induces a complete bipartite graph between $\beta(u)$ and $\beta(v)$ in $G$. Because we have to avoid double-counting, we obtain the following equation:
$\delta_{j\ell}=\sum_{u\in T_{j}}\sum_{\begin{subarray}{c}v\in T_{\ell},\\ uv\in B(G_{i-1})\end{subarray}}f_{j}(u)\cdot|\beta(v)|+f_{\ell}(v)\cdot|\beta(u)|-f_{j}(u)\cdot f_{\ell}(v)$
Using this equation, we can compute an optimal solution of $\pi$. The rest of the proof of Theorem 3.17 is analogous to the proof of Theorem 3.15.
The Details for Partial Vertex Cover
In this section, we show that Algorithm 2 solves the Partial Vertex Cover problem, thereby proving Theorem 3.17. We will reuse the notion of extended profiles, see Definition 3.1, and the notion of solutions, see Definition 3.2. However, we will not need the third component of a profile, and so it will always be set to $\emptyset$, i.e., a profile will be a tuple $(T,D,\emptyset,f)$. Instead of the vertex-cover-value of a solution, we will say simply the value. We say that a solution is an optimal solution of $\pi$ if no other solution of $\pi$ has higher value than $S$.
Input: A graph $G$, contraction sequence $(G_{1}=G,\ldots,G_{n})$ of width $d$, and $k\in\mathbb{N}$.
1 for each extended $1$-profile $\pi=(T,D,\emptyset,f)$ do  $\sigma_{1}(\pi):=(D,0)$
2 for $i$ from $2$ to $n$ do
3       $v:=$ the new vertex in $G_{i}$
4       $u_{1},u_{2}:=$ the two vertices of $G_{i-1}$ contracted into $v$
5       for each extended $i$-profile $\pi=(T,D,\emptyset,f)$ do
6             if $v\notin T$ then $\sigma_{i}(\pi):=\sigma_{i-1}(\pi)$; continue with next $\pi$
7             $\sigma_{i}(\pi):=(\textsf{null},-1)$
8             $T^{\prime}:=(T\setminus v)\cup\{u_{1},u_{2}\}$
9             for each function $f^{\prime}\colon T^{\prime}\rightarrow[0,k]$ compatible with $\pi$ do
10                   Let $\{\pi_{j}\colon j\in[m]\}$ be an $f^{\prime}$-decomposition of $\pi$ s.t. $m\leq d+2$; it exists by Lemma 3.5.
11                   for each $j\in[m]$ do
12                         $(T_{j},D_{j},\emptyset,f_{j}):=\pi_{j}$; $(S_{j},\nu_{j}):=\sigma_{i-1}(\pi_{j})$
13
14                  for each $1\leq j<\ell\leq m$ do
15                         $\textsf{value}_{j\ell}:=0$
16                         for each $x\in T_{j}$, $y\in T_{\ell}$ s.t. $xy\in B(G_{i-1})$ do
17                               $\textsf{value}_{j\ell}:=\textsf{value}_{j\ell}+f_{j}(x)\cdot|\beta(y)|+f_{\ell}(y)\cdot|\beta(x)|-f_{j}(x)\cdot f_{\ell}(y)$
18
19
20                  $S:=\bigcup_{j\in[m]}S_{j}$; $\nu:=\sum_{j\in[m]}\nu_{j}+\sum_{1\leq j<\ell\leq m}\textsf{value}_{j\ell}$
21                   if $\nu\geq\nu_{0}$, where $(S_{0},\nu_{0})=\sigma_{i}(\pi)$ then $\sigma_{i}(\pi):=(S,\nu)$
22
23
24
25$(S,\nu):=\sigma_{n}(\{u\},\{u\},\{u\},\{(u,k)\})$, where $\{u\}:=V(G_{n})$
return $\nu\geq t$
Algorithm 2 Algorithm for Partial Vertex Cover
We will proceed analogously to Section 3.3. The following lemma is the counterpart of Lemma 3.11.
Lemma 3.18.
Let $i\in[2,n]$, $\pi=(T,D,\emptyset,f)$ be an extended $i$-profile, $f^{\prime}$ be a function compatible with $\pi$, and $\nu$ and $S$ be the values computed by Algorithm 2 on line 2 in the iteration processing $f^{\prime}$. If for each extended $(i-1)$-profile $\pi^{\prime}$, $S^{\prime}$ is a solution of $\pi^{\prime}$ of value $\nu^{\prime}$, where $(S^{\prime},\nu^{\prime})$ is the pair computed in $\sigma_{i-1}(\pi^{\prime})$, then $S$ is a solution of $\pi$ of value $\nu$.
Proof 3.19.
Let $\{\pi_{j}\colon j\in[m]\}$ be the $f^{\prime}$-decomposition of $\pi$ found on line 2 in the iteration processing $f^{\prime}$, and for each $j\in[m]$, let $T_{j},D_{j},f_{j},S_{j}$, and $\nu_{j}$ be the values computed on line 2. By Lemma 3.7, $S$ is a solution of $\pi$. To simplify notation, we denote $\delta(A,B)=|\{uv\in E(G)\colon u\in\beta(A),v\in\beta(B),\{u,v\}\cap S\neq\emptyset\}|$ for any $A,B\subseteq V(G_{i})$ or $A,B\subseteq V(G_{i-1})$. Hence, by Definition 3.2, we need to show that $\nu=\delta(T,T)$.
Observe that $\delta(T,T)=\sum_{1\leq j\leq\ell\leq m}\delta(T_{j},T_{\ell})$. Since $\nu_{j}$ is the value of $S_{j}$ and $S\cap T_{j}=S_{j}$, $\nu_{j}=\delta(T_{j},T_{j})$ for any $j\in[m]$. Let us fix $j,\ell\in[m]$ such that $j<\ell$. Now it suffices to show that $\textsf{value}_{j\ell}=\delta(T_{j},T_{\ell})$ is true on line 2. Observe that $\delta(T_{j},T_{\ell})=\sum_{u\in T_{j},v\in T_{\ell}}\delta(\{u\},\{v\})$. Let us fix $u\in T_{j}$ and $v\in T_{\ell}$, and let $\delta(u,v):=\delta(\{u\},\{v\})$. Suppose that $\delta(u,v)>0$, which means that $u\in D_{j}$ or $v\in D_{\ell}$ because $S\cap\beta(\{u,v\})\neq\emptyset$, see Definition 3.1. Observe that $uv\notin R(G_{i-1})$ by condition 3 in the definition of a basic decomposition, see Section 2. Since $\delta(u,v)>0$, we obtain $uv\in B(G_{i-1})$. Finally, observe that in the iteration of the loop on lines 2-2, $\delta(u,v)$ is added to $\textsf{value}_{j\ell}$. Indeed, there are all possible edges between $\beta(u)$ and $\beta(v)$ in $G$, $f_{j}(u)=|S\cap\beta(u)|$ and $f_{\ell}(v)=|S\cap\beta(v)|$ by Definition 3.2, and the last term avoids double-counting the edges with both endpoints in $S$. We have shown that $\nu=\delta(T,T)$ as desired.
The following lemma is the counterpart of Lemma 3.13 in the partial dominating set algorithm.
Lemma 3.20.
For each $i\in[n]$ and each extended $i$-profile $\pi$, if the final value of $\sigma_{i}(\pi)$ computed by Algorithm 2 is $(S,\nu)$, then $S$ is an optimal solution of $\pi$ and $\nu$ is the value of $S$.
Proof 3.21.
Let $i\in[n]$ and $\pi=(T,D,\emptyset,f)$ be an extended $i$-profile. First, suppose that $i=1$. Since $T=\{u\}$ for some $u\in V(G)$, there are no edges in $G[\beta(T)]=G[T]$, and the value of the only solution $D$ of $\pi$ is clearly 0, see line 2. Now suppose that $i>1$. Let $v\in V(G_{i})$ be the new vertex in $G_{i}$ and let $u_{1},u_{2}\in V(G_{i-1})$ be the two vertices contracted into $v$. If $v\notin T$, then $\pi$ is also an $(i-1)$-profile and, by induction hypothesis, $\sigma_{i}(\pi)=\sigma_{i-1}(\pi)$ is an optimal solution of $\pi$, see line 2. Hence, let us assume that $v\in T$ and let $T^{\prime}=(T\setminus v)\cup\{u_{1},u_{2}\}$. By induction hypothesis and Lemma 3.18, $\pi$ has a solution since there is a function compatible with $\pi$.
Let $S^{*}$ be an optimal solution of $\pi$, let $\nu^{*}$ be the value of $S^{*}$, and let $f^{\prime}\colon T^{\prime}\rightarrow[0,k]$ be defined as $f^{\prime}(u)=|S^{*}\cap\beta(u)|$. As in the proof of Lemma 3.13, $f^{\prime}$ is compatible with $\pi$. Let us consider the iteration of the loop on lines 2-2 processing $\pi$ and the iteration of the loop on lines 2-2 processing $f^{\prime}$. Let $\{\pi_{j}\colon j\in[m]\}$ be the $f^{\prime}$-decomposition of $\pi$ found on line 2. For each $j\in[m]$, let $T_{j},D_{j},f_{j},S_{j}$ and $\nu_{j}$ be as in the algorithm. Let $\nu$ and $S$ be the values computed on line 2. By Lemma 3.18, $S$ is a solution of $\pi$ with value $\nu$. Similarly to the proof of Lemma 3.18, we denote $\delta(A,B)=|\{uv\in E(G)\colon u\in\beta(A),v\in\beta(B),\{u,v\}\cap S\neq\emptyset\}|$ and $\delta^{*}(A)=|\{uv\in E(G)\colon u\in\beta(A),v\in\beta(B),\{u,v\}\cap S^{*}\neq\emptyset\}|$ for any $A\subseteq V(G_{i})$ or $A\subseteq V(G_{i-1})$.
Suppose for contradiction that $\nu<\nu^{*}$, i.e., $\delta(T,T)<\delta^{*}(T,T)$. Observe that $\delta(T,T)=\sum_{1\leq j\leq\ell\leq m}\delta(T_{j},T_{\ell})$ and $\delta^{*}(T,T)=\sum_{1\leq j\leq\ell\leq m}\delta^{*}(T_{j},T_{\ell})$. Hence, there are $j,\ell\in[m]$ such that $\delta(T_{j},T_{\ell})<\delta^{*}(T_{j},T_{\ell})$. Observe that for each $j,\ell\in[m]$, $j\neq\ell$, we have
$\delta(T_{j},T_{\ell})=\sum_{\begin{subarray}{c}u\in T_{j},v\in T_{\ell},\\ uv\in B(G_{i-1})\end{subarray}}f_{j}(u)\cdot|\beta(v)|+f_{\ell}(v)\cdot|\beta(u)|-f_{j}(u)\cdot f_{\ell}(v)=\delta^{*}(T_{j},T_{\ell}).$
Indeed, if $u\in T_{j},v\in T_{\ell}$, and $uv\in R(G_{i-1})$, then $f_{j}(u)=0=f_{\ell}(v)$ by condition 3 in the definition of a basic decomposition, see Section 2. Hence, there is $j\in[m]$ such that $\delta(T_{j},T_{j})<\delta^{*}(T_{j},T_{j})$. However, $\delta(T_{j},T_{j})$ is the value of $S_{j}$ and $\delta^{*}(T_{j},T_{j})$ is the value of $S^{*}\cap\beta(T_{j})$, which is a solution of $\pi_{j}$. This is a contradiction since $S_{j}$ is an optimal solution of $\pi_{j}$ by induction hypothesis. Hence, $\nu=\nu^{*}$, and $S$ is an optimal solution of $\pi$, too.
After the iteration of the loop on lines 2-2 processing $f^{\prime}$, $\sigma_{i}(\pi)=(S,\nu)$. Observe that the second component of $\sigma_{i}(\pi)$ cannot change in any of the following iterations because $\nu$ is always the value of some solution of $\pi$ by Lemma 3.18. Hence, after all functions compatible with $\pi$ have been processed, $\sigma_{i}(\pi)$ is indeed an optimal solution of $\pi$ and its value, as desired.
We are now ready to prove the main theorem of this section.
Proof 3.22 (Proof of Theorem 3.17).
Let $\iota=(G,C=(G=G_{1},\ldots,G_{n}),k,t)$ be the input instance, where $C$ is contraction sequence of width $d$. Let $V(G_{n})=\{u\}$ and let $\pi=(\{u\},\{u\},\emptyset,\{(u,k)\})$. Note that $\pi$ is an extended $n$-profile unless $n<k$, in which case, $(G,k,t)$ is a NO instance. By Lemma 3.20, $S$ is an optimal solution of $\pi$ with value $\nu$, where $(S,\nu)$ is the pair computed by Algorithm 2. By Definition 3.2, $|S|=k$ and $\nu=|\{uv\in E(G)\colon u\in S,v\in\beta(\{u\})\}|=|\{uv\in E(G)\colon u\in S\}|$. Hence, $(G,k,t)$ is a YES instance if and only if $\nu\geq t$.
The running time of Algorithm 2 follows from the same arguments given in the proof of Theorem 3.15 used for Algorithm 1.
4 ($\exists^{*}\#$QF) Model Checking
We now turn our attention to the framework mentioned in the introduction. We consider a fragment of the counting logic FO(>0), namely problems expressible by formulas of the form $\exists x_{1}\dots\exists x_{k}\#y\,\psi(x_{1},\dots x_{k},y)\geq t$ for some quantifier-free formula $\psi$ and $t\in\mathbb{N}$. This fragment was also considered in [17]. For a background in (counting) logic, we refer the readers to [30]. Here, we want to mention that the counting quantifier #$\#y\,\psi(v_{1},\dots,v_{k},y)\geq t$ (for a given FO-formula $\psi$, vertices $v_{1},\dots,v_{k}$ assigned to variables $x_{1},\ldots,x_{k}$, and $t\in\mathbb{N}$) is true if there exist at least $t$ assignments to $y$ such that $\psi$ is satisfied. In our case, the variables always range over the set of vertices.
We are interested to give an fpt algorithm on graph classes of bounded twin-width to solve the model-checking problem for the fragment mentioned above.
($\exists^{*}\#$QF) Model Checking using twin-width
Input: A graph $G$, a contraction sequence for $G$ of width $d$, a quantifier-free formula $\psi$ with $k+1$ free variables, and $t\in\mathbb{N}$
Question: Are there vertices $v_{1},\dots,v_{k}\in V(G)$ such that $G\models\#y\,\psi(v_{1}\dots v_{k},y)\geq t$?
Parameter: $k+|\psi|+d$
Example 4.1.
We can express Partial Dominating Set as an instance of the Model-Checking problem above. Let $\psi(x_{1},\dots,x_{k},y)$ express that $y$ is dominated by a vertex from $x_{1},\dots,x_{k}$. That is, $\psi(x_{1},\dots,x_{k},y):=\bigvee_{i\in[k]}E(x_{i},y)\lor x_{i}=y$. Then, vertices $v_{1},\dots,v_{k}$ form a partial dominating set dominating $t$ vertices if and only if $G\models\#y\,\psi(v_{1},\dots v_{k},y)\geq t$.
Note that the Partial Vertex Cover cannot be expressed in this logic; the counting quantifier # is not allowed to range over edges. Also, the natural construction, where the edges of the input graph are subdivided, and the formula expresses the partial dominating set problem on the subdivision vertices, does not maintain bounded twin-width [4].
4.1 Virtual Profiles
In Section 3, we used extended profiles, which were obtained by “extending” basic profiles introduced in Section 2. In our model checking algorithm, we will again “extend” the basic profiles in a certain way. This time, however, these profiles will need to contain even more information than the extended profiles. Intuitively, the reason behind this change is that the $k$ existential variables in the considered formula are not interchangeable, whereas the $k$ vertices in a partial dominating set (or vertex cover) are. Hence, a part of a virtual profile is a function $f$, which describes where each existential variable should be realized.
However, we need to be able to capture the fact that some variables may be realized in a part of $G$ not covered by the profile; this is done by adding a bounded number of “virtual” vertices and edges ($\mathcal{V}$ and $\mathcal{E}$). Note that the virtual vertices are “new objects”; e.g., $V(G)\cap\mathcal{V}=\emptyset$. Moreover, we need to remember which atomic formulas ($x_{a}=x_{b}$, $E(x_{a},x_{b})$, and negations of these two) are satisfied by which existential variables: this information is encoded in a vertex-labeled graph $H$. Note that if a virtual profile has a solution, then $H$ must be compatible with $f$ and $\mathcal{E}$.
Definition 4.2.
Let $i\in[n]$ and let $(T,D)$ be a basic $i$-profile. A virtual $i$-profile is a tuple $\pi=(T,D,\mathcal{V},\mathcal{E},f,H)$, where:
  • $\mathcal{V}$ is the set of at most $k$ virtual vertices.
  • $\mathcal{E}\subseteq\{uv\colon u\in\mathcal{V}\cup T,v\in\mathcal{V}\}$ is the set of virtual edges.
  • $f\colon[k]\rightarrow D\cup\mathcal{V}$ is a surjective function such that $|f^{-1}(u)|\leq|\beta(u)|$ for each $u\in D$.
  • $H$ is a graph with at most $k$ vertices with vertex set labeled as $\{h_{1},\ldots,h_{k}\}$.
The expanded graph of $\pi$, denoted $G_{\pi}$, is the graph with vertex set $\beta(T)\cup\mathcal{V}$ and edge set $E(G[\beta(T)])\cup\{uv\in\mathcal{E}\colon u,v\in\mathcal{V}\}\cup\{u_{0}v\colon uv\in\mathcal{E},v\in\mathcal{V},u\in T,u_{0}\in\beta(u)\}$. A tuple $\mathbf{s}=(s_{1},\ldots,s_{k})\in V(G_{\pi})^{k}$ is a solution of $\pi$ if:
  • for each $a\in[k]$, $s_{a}=f(a)$ if $f(a)\in\mathcal{V}$, and $s_{a}\in\beta(f(a))$ if $f(a)\in D$;
  • the function $s_{a}\mapsto h_{a}$ is an isomorphism from $G_{\pi}[\{s_{1},\ldots,s_{k}\}]$ to $H$.
The value of $\mathbf{s}$ is defined as |$|\{u\in\beta(T)\colon G_{\pi}\models\psi(s_{1},\ldots,s_{k},u)\}|$. A solution $\mathbf{s}$ of $\pi$ is optimal if no other solution of $\pi$ has higher value than $\mathbf{s}$.
The expanded graph $G_{\pi}$ illustrates the concept of virtual profiles: the subgraph of $G$ induced by the bags from $T$ is expanded with the virtual vertices. The virtual vertices are uniformly connected to vertices from each bag in $T$ as dictated by $\mathcal{E}$. A solution $\mathbf{s}$ of $\pi$ must then adhere to $\pi$ in the following way: each $s_{a}$ in $\mathbf{s}$ has to be the correct virtual vertex or a vertex from the correct bag as dictated by $f$; and $\mathbf{s}$ has to induce the desired subgraph $H$ in the expanded graph $G_{\pi}$. When an optimal solution has been computed for each virtual profile, we can find the answer to the problem by inspecting an $n$-profile $\pi^{*}$ such that $\mathcal{V}$ and $\mathcal{E}$ are empty, $T=V(G_{n})$ and $\beta(T)=V(G)$, and $f$ is the unique function from $[k]$ to $V(G_{n})$. In contrast to the case of Partial Dominating Set, the profile $\pi^{*}$ is not unique; one has to try all possible choices of $H$.
Now we show that the number of virtual profiles is bounded.
Lemma 4.3.
If $i\in[n]$ and $v\in V(G_{i})$, then the number of virtual $i$-profiles $(T,D,\mathcal{V},\mathcal{E},f,H)$ such that $v\in T$ is in $2^{\mathcal{O}(k^{2}d\log(d))}$. Moreover, such profiles can be enumerated in time $2^{\mathcal{O}(k^{2}d\log(d))}$.
Proof 4.4.
By Lemma 2.4, there is $d^{\mathcal{O}(kd)}$ basic $i$-profiles $(T,D)$. Let us bound the number of possible graphs $H$. There are at most $k\cdot 2^{k^{2}}$ graphs on at most $k$ vertices and for each of them, there are at most $k^{k}$ vertex-labelings with labels $\{h_{1},\ldots,h_{k}\}$. Hence, there are $2^{\mathcal{O}(k^{2})}$ choices of $H$. There are $k+1$ choices of $\mathcal{V}$ (it suffices to choose the size of $\mathcal{V}$) and at most $2^{k\cdot(k+k(d+1))}\in 2^{\mathcal{O}(k^{2}d)}$ choices of $\mathcal{E}$. Finally, there are at most $(k+k(d+1))^{k}\in 2^{\mathcal{O}(k\log(kd))}$ choices of $f$. Hence, there are $2^{\mathcal{O}(k^{2}d\log(d))}$ virtual $i$-profiles containing $v$, and the claim about enumeration follows from Lemma 2.4.
4.2 Virtual Decompositions
To achieve our goal of finding optimal solutions of virtual profiles, we continue similarly as in the previous section about Partial Dominating Set and Partial Vertex Cover. We will again need a notion of a decomposition for our virtual profiles. However, we must ensure that the profiles in the decomposition are “compatible” with each other with respect to the virtual vertices. Informally, the virtual edges must correspond to the black edges between the profiles in $G_{i-1}$.
Let $i\in[2,n]$, let $v\in V(G_{i})$ be the new vertex in $G_{i}$, let $u_{1},u_{2}\in V(G_{i-1})$ be the two vertices contracted into $v$, let $\pi_{0}=(T,D)$ be a basic $i$-profile such that $v\in T$, let $\pi=(T,D,\mathcal{V},\mathcal{E},f,H)$ be a virtual $i$-profile, let $T^{\prime}=T\setminus\{v\}\cup\{u_{1},u_{2}\}$, and let =$\mathcal{E}^{\prime}=\{uw\in\mathcal{E}\colon u,w\in T^{\prime}\cup\mathcal{V}\}\cup\{u_{1}w,u_{2}w\colon vw\in\mathcal{E}\}$. We say that a function $f^{\prime}\colon[k]\rightarrow T^{\prime}\cup\mathcal{V}$ is compatible with $\pi$ if $f^{\prime}(a)\in\{u_{1},u_{2}\}$ when $f(a)=v$, and $f^{\prime}(a)=f(a)$ otherwise.
111133335656565644442222
111123423423423456565656
111122223434343456565656
222211113333565656564444
343434345656565611112222
Figure 3: Leftmost: A graph $H$ with vertex set $\{h_{1},\ldots,h_{6}\}$. The circle representing $h_{a}$ contains the digit $a$. Middle left: A virtual $i$-profile $\pi=(T,D,\mathcal{V},\mathcal{E},f,H)$ for $k=6$. Virtual vertices are represented by violet stars and virtual edges by violet lines. $T$ is the set of all non-virtual vertices. For each $a\in[6]$, the vertex $f(a)$ contains the digit $a$. $D$ is the set $f(\{2,5\})$. Middle: The circles, and black and red edges represent $G_{i-1}[T^{\prime}]$, the violet star is the vertex in $\mathcal{V}$, and the violet lines represent $\mathcal{E}^{\prime}$. The digits represent the function $f^{\prime}$. The set $T^{\prime}$ has two red-connected components: $T_{1}$ colored in green and $T_{2}$ colored in orange. Middle right: The profile $(T_{1},D_{1},\mathcal{V}_{1},\mathcal{E}_{1},f_{1},H)$. Note that the facts that $f_{1}(5)=f_{1}(6)$ and $f_{1}(3)f_{1}(4)\notin\mathcal{E}_{1}$ follow from $H$; it cannot be deduced from $T$ or $T^{\prime}$. Rightmost: The profile $(T_{2},D_{2},\mathcal{V}_{2},\mathcal{E}_{2},f_{2},H)$.
If a function $f^{\prime}$ is compatible with $\pi$, then a set $\{\pi_{j}=(T_{j},D_{j},\mathcal{V}_{j},\mathcal{E}_{j},f_{j},H)\colon j\in[m]\}$ of virtual $(i-1)$-profiles is an $f^{\prime}$-decomposition if $D^{\prime}=\operatorname{range}(f^{\prime})\cap T^{\prime}$, $\{(T_{1},D_{1}),\ldots,(T_{m},D_{m})\}$ is a $D^{\prime}$-decomposition of $\pi_{0}$, and for each $j\in[m]$:
  • $\mathcal{V}_{j}=\mathcal{V}\cup\{h_{a}\colon f^{\prime}(a)\in D^{\prime}\setminus D_{j}\}$.
  • $f_{j}(a)=f^{\prime}(a)$ if $f^{\prime}(a)\in D_{j}\cup\mathcal{V}$, and $f_{j}(a)=h_{a}$ if $f^{\prime}(a)\in D^{\prime}\setminus D_{j}$.
  • $\mathcal{E}_{j}$ is the union of the following sets.
    • $\{uw\in\mathcal{E}^{\prime}\colon u,w\in\mathcal{V}\cup T_{j}\}$.
    • $\{uh_{a}\colon u\in\mathcal{V}\cup T_{j},f^{\prime}(a)=w\in D^{\prime}\setminus D_{j},uw\in\mathcal{E}^{\prime}\cup B(G_{i-1})\}$.
    • $\{h_{a}h_{b}\in E(H)\colon f^{\prime}(a),f^{\prime}(b)\in D^{\prime}\setminus D_{j}\}$.
The intuition behind this construction of $\pi_{j}$ is that for each vertex in $D^{\prime}\setminus D_{j}$, which are the vertices with some variable assigned to them, we add a new virtual vertex to $\mathcal{V}_{j}$, and $\mathcal{E}_{j}$ contains $\mathcal{E}^{\prime}$ “restricted” to $T_{j}$ (the first term), edges between a new virtual vertex $h_{a}$ and an “old” virtual vertex $u\in\mathcal{V}$ or a vertex $u\in T_{j}$ (the second term), and edges between two new virtual vertices (the third term). Note that we used the vertices of $H$ to label the new virtual vertices, which means that two new virtual vertices $h_{a}$ and $h_{b}$ may be equal even when $a\neq b$; this ensures the possibility of an isomorphism from a solution of $\pi_{j}$ to $H$. See Figure 3 for an illustration of a virtual decomposition.
Again, we prove that a small $f^{\prime}$-decomposition exists for each $f^{\prime}$ compatible with $\pi$.
Lemma 4.5.
If $i\in[2,n]$, $V(G_{i})\setminus V(G_{i-1})=\{v\}$, $V(G_{i-1})\setminus V(G_{i})=\{u_{1},u_{2}\}$, $\pi=(T,D,\mathcal{V},\mathcal{E},f,H)$ is a virtual $i$-profile such that $v\in T$, $T^{\prime}=T\setminus\{v\}\cup\{u_{1},u_{2}\}$, and $f^{\prime}\colon[k]\rightarrow T^{\prime}\cup\mathcal{V}$ is a function compatible with $\pi$, then there is an $f^{\prime}$-decomposition $\mathcal{D}$ of $\pi$ with cardinality at most $d+2$. Moreover, $\mathcal{D}$ can be found in time $\mathcal{O}(k^{2}d^{2})$.
Proof 4.6.
Let $D^{\prime}=\operatorname{range}(f^{\prime})\cap T^{\prime}$. By Lemma 2.2, a $D^{\prime}$-decomposition $\mathcal{D}_{0}=\{(T_{j},D_{j})\colon j\in[m]\}$ of $(T,D)$ such that $m\leq d+2$ can be found in time $\mathcal{O}(kd^{2})$. For each $j\in[m]$, define $\mathcal{V}_{j}$, $\mathcal{E}_{j}$, and $f_{j}$ as in the definition of an $f^{\prime}$-decomposition above. Now it can be easily seen that $\{(T_{j},D_{j},\mathcal{V}_{j},\mathcal{E}_{j},f_{j},H)\colon j\in[m]\}$ is the desired $f^{\prime}$-decomposition of $\pi$.
Let $j\in[m]$ and observe that computing $\mathcal{V}_{j}$ and $f_{j}$ is straightforward. To compute $\mathcal{E}_{j}$, we have to inspect all pairs $u\in T_{j}$ and $v\in D^{\prime}\setminus D_{j}$. The number of all such pairs, for all $j\in[m]$ in total, is bounded by $|T^{\prime}|\cdot|D^{\prime}|\in\mathcal{O}(k^{2}d)$. After adding the time required to find $\mathcal{D}_{0}$, we obtain the stated running time.
4.3 Combining Solutions
In Section 3, we obtained a solution for an extended $i$-profile $\pi$ by simply taking the union of solutions of profiles in a decomposition of $\pi$. In this section, combining solutions is more complicated because each solution is a tuple, which may contain some virtual vertices.
Definition 4.7.
Let $i\in[n]$, let $\pi_{1}=(T_{1},D_{1},\mathcal{V}_{1},\mathcal{E}_{1},f_{1},H)$ and $\pi_{2}=(T_{2},D_{2},\mathcal{V}_{2},\mathcal{E}_{2},f_{2},H)$ be two virtual $i$-profiles, and let $\mathbf{s}^{1}=(s_{1}^{1},\ldots,s_{k}^{1})$ and $\mathbf{s}_{2}=(s_{1}^{2},\ldots,s_{k}^{2})$ be solutions of $\pi_{1}$ and $\pi_{2}$, respectively. By $\mathbf{s}^{1}\oplus\mathbf{s}^{2}$, we denote the tuple $(s_{1},\ldots,s_{k})\in(\beta(D_{1})\cup V(G_{\pi_{2}}))^{k}$ such that for each $a\in[k]$, $s_{a}=s_{a}^{1}$ if $f_{1}(a)\in D_{1}$, and $s_{a}=s_{a}^{2}$ otherwise.
Note that in general, the operation $\oplus$ is neither associative nor commutative. However, we will use it only to combine solutions of profiles in a decomposition (and when restricted to such inputs, $\oplus$ is both associative and commutative, which will be implicit in the proof of the following lemma). Let $\mathbf{s}=\bigoplus_{j\in[m]}\mathbf{s}^{j}=\mathbf{s}^{1}\oplus(\mathbf{s}^{2}\oplus\ldots)$. Formally, $\mathbf{s}=\mathbf{s}^{1}$ if $m=1$ and $\mathbf{s}=\mathbf{s}^{1}\oplus\bigoplus_{j\in[2,m]}\mathbf{s}^{j}$ otherwise.
Lemma 4.8.
If $i\in[2,n]$, $\pi$ is a virtual $i$-profile, $\{\pi_{j}\colon j\in[m]\}$ is an $f^{\prime}$-decomposition of $\pi$, and $\mathbf{s}^{j}$ is a solution of $\pi_{j}$ for each $j\in[m]$, then $\mathbf{s}=\bigoplus_{j\in[m]}\mathbf{s}^{j}$ is a solution of $\pi$. Moreover, if $\nu_{j}$ is the value of $\mathbf{s}_{j}$, then the value of $\mathbf{s}$ is $\sum_{j\in[m]}\nu_{j}$.
Proof 4.9.
Let $\pi=(T,D,\mathcal{V},\mathcal{E},f,H)$ and $\mathbf{s}=(s_{1},\ldots,s_{k})$. For each $j\in[m]$, let $\pi_{j}=(T_{j},D_{j},\mathcal{V}_{j},\mathcal{E}_{j},f_{j},H)$ and let $\mathbf{s}_{j}=(s_{1}^{j},\ldots,s_{k}^{j})$. For each $j\in[m]$, let $G^{j}:=G_{\pi_{j}}$.
First, let $a\in[k]$ be such that $f(a)\in\mathcal{V}$. Observe that for each $j\in[m]$, $f(a)=f^{\prime}(a)=f_{j}(a)=s_{a}^{j}$, which means that $s_{a}=f(a)$ as desired (because $s_{a}=s_{a}^{j}$ for some $j\in[m]$ by Definition 4.7). Second, let $a\in[k]$ be such that $f(a)\in D$ and let $j\in[m]$ be the unique index such that $u:=f^{\prime}(a)=f_{j}(a)\in D_{j}$. By Definition 4.7, $s_{a}=s_{a}^{j}\in\beta(u)$ because $f_{\ell}(a)\notin D_{\ell}$ for each $\ell\in[m]\setminus\{j\}$. By definition of $f^{\prime}$, if $u\in V(G_{i-1})\cap V(G_{i})$, then $f(a)=u$. Otherwise, $f(a)$ is the new vertex in $G_{i}$, and $s_{a}\in\beta(u)\subseteq\beta(f(a))$. In both cases, $s_{a}\in\beta(f(a))$ as desired.
Now we need to show that the function $s_{a}\mapsto h_{a}$ is an isomorphism from $G_{\pi}[\{s_{1},\ldots,s_{k}\}]$ to $H$. Let us fix $a,b\in[k]$. Since $\mathbf{s}_{j}$ is a solution of $\pi_{j}$ for each $j\in[m]$, it suffices to show that $s_{a}s_{b}\in E(G_{\pi})\leftrightarrow s_{a}^{j}s_{b}^{j}\in E(G^{j})$ and $s_{a}=s_{b}\leftrightarrow s_{a}^{j}=s_{b}^{j}$ for some $j\in[m]$. First, if $s_{a},s_{b}\in\mathcal{V}$ or $s_{a},s_{b}\in\beta(T_{j})$ for some $j\in[m]$, then $s_{a}=s_{a}^{j}$ and $s_{b}=s_{b}^{j}$, and $s_{a}s_{b}\in E(G_{\pi})\leftrightarrow s_{a}s_{b}\in E(G)\cup\mathcal{E}\leftrightarrow s_{a}^{j}s_{b}^{j}\in E(G^{j})$ as required. Moreover, we clearly have $s_{a}=s_{b}\leftrightarrow s_{a}^{j}=s_{b}^{j}$; note that in all other cases, $s_{a}\neq s_{b}$ and $s_{a}^{j}\neq s_{b}^{j}$. Second, if $s_{a}\in\mathcal{V}$ and $s_{b}\in\beta(u)\subseteq\beta(u^{\prime})$ for some $u\in T_{j}$, $j\in[m]$, and $u^{\prime}\in T$, then $s_{a}s_{b}\in E(G_{\pi})\leftrightarrow s_{a}u^{\prime}\in\mathcal{E}\leftrightarrow s_{a}u\in\mathcal{E}_{j}\leftrightarrow s_{a}s_{b}\in E(G^{j})$. Finally, assume that $j,\ell\in[m]$, $j\neq\ell$, and $s_{a}\in\beta(T_{j}),s_{b}\in\beta(T_{\ell})$. Let $u=f^{\prime}(a)=f_{j}(a)$ and $v=f^{\prime}(b)=f_{\ell}(b)$. By definition of a solution, we have $s_{a}\in\beta(u)$ and $s_{b}\in\beta(v)$. Recall that $D^{\prime}=\bigcup_{p\in[m]}D_{p}$, that $u,v\in D^{\prime}$, and that $\{(T_{p},D_{p})\colon p\in[m]\}$ is a $D^{\prime}$-decomposition of the basic $i$-profile $(T,D)$. Hence, by condition 3 in the definition of a $D^{\prime}$-decomposition, see Section 2, $uv\notin R(G_{i-1})$. Now $s_{a}s_{b}\in E(G_{\pi})\leftrightarrow s_{a}s_{b}\in E(G)\leftrightarrow uv\in B(G_{i-1})\leftrightarrow uh_{b}\in\mathcal{E}_{j}\leftrightarrow s_{a}^{j}s_{b}^{j}\in E(G^{j})$ as required. Note that after the third equivalence, we could have used $h_{a}v\in\mathcal{E}_{\ell}$ and continued analogously. We have shown that $\mathbf{s}$ is a solution of $\pi$.
What remains to be shown is that the value of $\mathbf{s}$ equals $\sum_{j\in[m]}\nu_{j}$. Since $\beta(T)=\beta(T^{\prime})$, where $T^{\prime}=\bigcup_{j\in[m]}T_{j}$, it suffices to show that for each $j\in[m]$ and $u\in\beta(T_{j})$, $G_{\pi}\models\psi(s_{1},\ldots,s_{k},u)$ if and only if $G^{j}\models\psi(s_{1}^{j},\ldots,s_{k}^{j},u)$. Using the isomorphisms to $H$, we obtain that $s_{a}\mapsto s_{a}^{j}$ is an isomorphism from $G_{\pi}[\{s_{1},\ldots,s_{k}\}]$ to $G^{j}[\{s_{1}^{j},\ldots,s_{k}^{j}\}]$. Hence, it suffices to show that for each $a\in[k]$, (1) $u=s_{a}\leftrightarrow u=s_{a}^{j}$, and (2) $us_{a}\in E(G_{\pi})\leftrightarrow us_{a}^{j}\in E(G^{j})$. If $s_{a}=s_{a}^{j}$, then (1) clearly holds, and if $s_{a}\neq s_{a}^{j}$, then $s_{a}\notin\beta(T_{j})$ and $s_{a}^{j}\in\mathcal{V}_{j}$, which implies $u\notin\{s_{a},s_{a}^{j}\}$. Now let us focus on (2). Let $u_{j}\in T_{j}$ and $u^{\prime}\in T$ be such that $u\in\beta(u_{j})\subseteq\beta(u^{\prime})$. If $s_{a}=s_{a}^{j}\in\beta(T_{j})$, then $us_{a}\in E(G_{\pi})\leftrightarrow us_{a}\in E(G)\leftrightarrow us_{a}\in E(G^{j})$ as required. If $s_{a}=s_{a}^{j}\notin\beta(T_{j})$, then $s_{a}\in\mathcal{V}$, and $us_{a}\in E(G_{\pi})\leftrightarrow u^{\prime}s_{a}\in\mathcal{E}\leftrightarrow u_{j}s_{a}\in\mathcal{E}_{j}\leftrightarrow us_{a}\in E(G^{j})$ as required. Finally, suppose $s_{a}\neq s_{a}^{j}$, i.e., $s_{a}\in\beta(T)\setminus\beta(T_{j})$, and $s_{a}^{j}=h_{a}\in\mathcal{V}_{j}$. Let $w=f^{\prime}(a)$ and observe that, by definition of a solution, $s_{a}\in\beta(w)$. Since $w\in D^{\prime}$, we have $u_{j}w\notin R(G_{i-1})$, see condition 3 in the definition of a $D^{\prime}$-decomposition. Hence, we have $us_{a}\in E(G_{\pi})\leftrightarrow us_{a}\in E(G)\leftrightarrow u_{j}w\in B(G_{i-1})\leftrightarrow u_{j}h_{a}\in\mathcal{E}_{j}\leftrightarrow uh_{a}=us_{a}^{j}\in E(G^{j})$ as required. We have proven that the value of $\mathbf{s}$ is $\sum_{j\in[m]}\nu_{j}$ as desired.
4.4 Model Checking Algorithm
Input: A graph $G$, contraction sequence $(G_{1}=G,\ldots,G_{n})$ of width $d$, and a formula $\psi(x_{1}\dots x_{k},y)$.
1
2for each virtual $1$-profile $\pi=(T,D,\mathcal{V},\mathcal{E},f,H)$ do
3       $\mathbf{s}:=(f(1),\ldots,f(k))$
4       if $\mathbf{s}$ is a solution of $\pi$ then $\sigma_{1}(\pi):=(\mathbf{s},|\{y\in T\colon G_{\pi}\models\psi(\mathbf{s},y)\}|)$
5       else $\sigma_{1}(\pi):=(\textsf{null},0)$
6
7 for $i$ from $2$ to $n$ do
8       $v:=$ the new vertex in $G_{i}$; $u_{1},u_{2}:=$ the two vertices of $G_{i-1}$ contracted into $v$
9       for each virtual $i$-profile $\pi=(T,D,\mathcal{V},\mathcal{E},f,H)$ do
10             if $v\notin T$ then $\sigma_{i}(\pi):=\sigma_{i-1}(\pi)$; continue with next $\pi$
11             $\sigma_{i}(\pi):=(\textsf{null},0)$
12             $T^{\prime}:=(T\setminus v)\cup\{u_{1},u_{2}\}$
13             for each function $f^{\prime}\colon[k]\rightarrow T^{\prime}\cup\mathcal{V}$ compatible with $\pi$ do
14                   Let $\{\pi_{j}\colon j\in[m]\}$ be an $f^{\prime}$-decomposition of $\pi$ s.t. $m\leq d+2$; it exists by Lemma 4.5.
15                   for each $j\in[m]$ do
16                        $(\mathbf{s}^{j},\nu_{j}):=\sigma_{i-1}(\pi_{j})$
17                         if $\mathbf{s}^{j}=\textsf{null}$ then continue with next $f^{\prime}$
18
19                  $\nu:=\sum_{j\in[m]}\nu_{j}$; $\mathbf{s}:=\bigoplus_{j\in[m]}\mathbf{s}^{j}$
20                   if $\nu\geq\nu_{0}$, where $\sigma_{i}(\pi):=(\mathbf{s}_{0},\nu_{0})$ then $\sigma_{i}(\pi):=(\mathbf{s},\nu)$
21
22
23
[…]
25 for each graph $H$ with vertex set labeled as $\{h_{1},\ldots,h_{k}\}$ do
26       $(\mathbf{s},\nu):=\sigma_{n}(\{u\},\{u\},\emptyset,\emptyset,f,H)$, where $f(a)=u$ for each $a\in[k]$
27       if $\nu\geq\textsf{count}$ then $\textsf{solution}:=\mathbf{s}$; $\textsf{count}:=\nu$
28
29 return $\textsf{count}\geq t$
Algorithm 3 Algorithm for ($\exists^{*}\#$QF) Model Checking
Now we can start proving the correctness of Algorithm 3. First, let us show that the algorithm computes an optimal solution for each virtual profile $\pi$ (or it discovers that $\pi$ has no solution).
Lemma 4.10.
For each $i\in[n]$ and each virtual $i$-profile $\pi$, if the final value of $\sigma_{i}(\pi)$ computed by Algorithm 3 is $(\mathbf{s},\nu)$, then $\mathbf{s}$ is an optimal solution of $\pi$ and $\nu$ is the value of $\mathbf{s}$. On the other hand, if $\sigma_{i}(\pi)=(\textsf{null},0)$, then $\pi$ has no solution.
Proof 4.11.
Let $i\in[n]$ and $\pi=(T,D,\mathcal{V},\mathcal{E},H)$ be a virtual $i$-profile. First, suppose that $i=1$. Since $G=G_{1}$ has no red edges, $T=\{u\}$ for some $u\in V(G)$. Recall that $\beta(u)=\{u\}$. Hence, by Definition 4.2, $(f(1),\ldots,f(k))$ is the only possible solution of $\pi$, and its value is either 0 or 1, see line 3. Now suppose that $i>1$. Let $v\in V(G_{i})$ be the new vertex in $G_{i}$ and let $u_{1},u_{2}\in V(G_{i-1})$ be the two vertices contracted into $v$. If $v\notin T$, then $\pi$ is also an $(i-1)$-profile and, by induction hypothesis, $\sigma_{i}(\pi)=\sigma_{i-1}(\pi)$ is an optimal solution of $\pi$ (or $(\textsf{null},0)$), see line 3. Hence, let us assume that $v\in T$ and let $T^{\prime}=(T\setminus v)\cup\{u_{1},u_{2}\}$.
If $\pi$ has no solution, then for each function $f^{\prime}$ compatible with $\pi$, the iteration of the loop on lines 3-3 processing $f^{\prime}$ must be exited on line 3; otherwise, a solution of $\pi$ would be found on line 3 (by induction hypothesis and Lemma 4.8). Hence, $\sigma_{i}(\pi)$ always equals $(\textsf{null},0)$ as required. Suppose that $\pi$ has a solution. Let $\mathbf{s}^{*}=(s_{1}^{*},\ldots,s_{k}^{*})$ be an optimal solution of $\pi$, let $\nu^{*}$ be the value of $S^{*}$, and let $f^{\prime}\colon[k]\rightarrow T^{\prime}\cup\mathcal{V}$ be such that $f^{\prime}(a)=s_{a}^{*}$ if $s_{a}^{*}\in\mathcal{V}$, and $s_{a}^{*}\in\beta(f^{\prime}(a))$ otherwise. Observe that $f^{\prime}$ is compatible with $\pi$, and let us consider the iteration of the loop on lines 3-3 processing $\pi$ and the iteration of the loop on lines 3-3 processing $f^{\prime}$. Let $\{\pi_{j}\colon j\in[m]\}$ be the $f^{\prime}$-decomposition of $\pi$ found on line 3. For each $j\in[m]$, let $\pi_{j}=(T_{j},D_{j},\mathcal{V}_{j},\mathcal{E}_{j},H)$, let $\mathbf{s}_{j}$ and $\nu_{j}$ be as in the algorithm, and let $\nu$ and $\mathbf{s}$ be the values computed on line 3. By Lemma 4.8, $\nu$ is the value of $\mathbf{s}$.
For each $j\in[m]$, let $\mathbf{s}^{j*}=(s_{1}^{j*},\ldots,s_{k}^{j*})$ be the solution of $\pi_{j}$ such that for each $a\in[k]$, $s_{a}^{j*}=s_{a}^{*}$ if $s_{a}^{*}\in\beta(T_{j})\cup\mathcal{V}$, and $s_{a}^{j*}=h_{a}$ otherwise. Observe that $\mathbf{s}^{*}=\bigoplus_{j\in[m]}\mathbf{s}^{j*}$, and so, by Lemma 4.8, $\nu^{*}=\sum_{j\in[m]}\nu^{*}_{j}$, where $\nu^{*}_{j}$ is the value of $\mathbf{s}^{j*}$. Suppose for contradiction that $\nu<\nu^{*}$. Since $\beta(T)=\beta(T^{\prime})$ and $T^{\prime}=\bigcup_{j\in[m]}T_{j}$, there is $j\in[m]$ such that $\nu_{j}<\nu^{*}_{j}$, which is a contradiction since $\mathbf{s}^{j}$ is an optimal solution of $\pi_{j}$ by the induction hypothesis. Hence, $\mathbf{s}$ is an optimal solution of $\pi$, too.
After the iteration of the loop on lines 3-3 processing $f^{\prime}$, $\sigma_{i}(\pi)=(\mathbf{s},\nu)$. Observe that the second component of $\sigma_{i}(\pi)$ cannot change in any of the following iterations because $\nu$ is always the value of some solution of $\pi$ by Lemma 4.8. Hence, after all functions compatible with $\pi$ have been processed, $\sigma_{i}(\pi)$ is indeed an optimal solution of $\pi$ and its value, as desired.
Using Lemma 4.10, it is easy to show the correctness of Algorithm 3.
Theorem 4.12.
Algorithm 3 decides the ($\exists^{*}\#$QF) Model Checking using twin-width problem in time $2^{\mathcal{O}(k^{2}d\log(d))}n$.
Proof 4.13.
Suppose that the algorithm is given a graph $G$, a contraction sequence $(G_{1}=G,\ldots,G_{n})$ of width $d$, and a formula $\phi\equiv\exists x_{1}\ldots\exists x_{k}\#y\,\psi(x_{1},\ldots,x_{k},y)\geq t$. Let us say that a graph $H$ with vertex set labeled as $\{h_{1},\ldots,h_{k}\}$ is a template graph. Let us consider an iteration of the loop one lines 3-3 processing a template graph $H$, let $\pi_{H}=(\{u\},\{u\},\emptyset,\emptyset,f,H)$ be as on line 3, and let $(\mathbf{s}_{H},\nu_{H})$ be the tuple computed on line 3. By Lemma 4.10, $\mathbf{s}_{H}$ is an optimal solution of $\pi_{H}$ with value $\nu_{H}$, or $(\mathbf{s}_{H},\nu_{H})=(\textsf{null},0)$ if $\pi_{H}$ has no solution. Since there are no virtual vertices in $\pi_{H}$, if $\pi_{H}$ has a solution, then $\mathbf{s}_{H}$ is a tuple in $V(G)^{k}$, and $\nu_{H}=|\{u\in V(G)\colon G\models\psi(\mathbf{s},u)\}|$.
[…]
Now need to argue that Algorithm 3 achieves the desired running time. As in the proof of Theorem 3.15, we maintain a single function $\sigma$ (instead of $\sigma_{i}$ for each $i\in[n]$). First, observe that by Lemma 4.3, there are at most $2^{\mathcal{O}(k^{2}d\log(d))}\cdot n$ virtual $1$-profiles. Hence, executing the for loop on lines 3-3 takes time at most $2^{\mathcal{O}(k^{2}d\log(d))}\cdot n$. Second, observe that there are $n-1$ iterations of the loop on lines 3-3; let us consider the iteration processing $i\in[2,n]$ and let $v$ be the new vertex in $G_{i}$. Let $\pi=(T,D,\mathcal{V},\mathcal{E},f,H)$ be a virtual $i$-profile. Since $\sigma(\pi)$ has already been computed if $v\notin T$, it suffices to consider only profiles such that $v\in T$; these profiles can be enumerated in time $2^{\mathcal{O}(k^{2}d\log(d))}$ by Lemma 4.3. There are at most $2^{k}$ functions $f^{\prime}$ compatible with $\pi$ and for each of them, we can find an $f^{\prime}$-decomposition of size at most $m\leq d+2$ in time $\mathcal{O}(k^{2}d^{2})$ by Lemma 4.5. Hence, the running time of one iteration of the loop is $2^{\mathcal{O}(kd\log(kd))}$. Finally, since there are $2^{\mathcal{O}(k^{2})}$ choices of $H$, executing the loop on lines 3-3 takes time $2^{\mathcal{O}(k^{2})}$. In total, the running time is $2^{\mathcal{O}(k^{2}d\log(d))}\cdot n$, as required.
5 Conclusion
We have demonstrated that several optimization problems requiring counting can be efficiently solved on graphs with small twin-width. This encompasses not only well-known problems such as Partial Vertex Cover and Partial Dominating Set, but also all problems expressible by a restricted fragment of first-order counting logic. It would be desirable to extend this result by providing a more comprehensive fragment that accommodates additional problems. For bounded expansion, we can utilize formulas of the form $\exists x_{1}\ldots\exists x_{k}\#y\,\phi(x_{1},\ldots,x_{k})$, where $\phi$ is an arbitrary first-order formula rather than a quantifier-free one, and even slightly more complex formulas [18, 33]. Can we achieve the same for graphs of bounded twin-width? Moreover, we can handle even more complicated formulas with nested counting quantifiers approximately [18]. Once again, does the same or a similar result hold true for bounded twin-width? It should be noted that the best result for nowhere-dense classes is basically the same as our result for twin-width [17].